I can row upstream at 2mph and downstream at 8mph. How far can I row to get back to starting point in 5 hours.
I know answer is 16 miles as in 4 hours I can row 8 miles upstream and in one hour I can return for total of 5 hours and 8 miles each way for total of 16 miles but what is correct equation. I am stumped as to how to do this algebraically? Thanks for your help. cpmurphy@sbcglobal.net

Answer by mananth(16949)

let the distance he goes be x
Time upstream + time downstream = 5 hours
speed upstream = 2 mph
speed downstream = 8 mph

x/2 + x/8 = 5
LCD = 8
4x+x=8*5
5x=40
x=8 miles
you can go 8 miles far to complete the return in 5 hours